OpenStudy (anonymous):

how do you simplify the following expression, and rewrite it in an equivalent form with positive exponents. -51x^4y/(17 x^2y^7)

OpenStudy (anonymous):

\[\frac{-51x^4y}{17 x^2y^7}\]

OpenStudy (anonymous):

yes that is the problem

OpenStudy (anonymous):

\[-\frac{51}{17} \times \frac{x^4}{x^2} \times \frac{y}{y^7} \]

OpenStudy (anonymous):

so you just do it individually

OpenStudy (anonymous):

It's much easier that way. You don't have to, you could write the fraction in any order that you want.